### 17 Mar 2026 v2

# Towards heterogeneous parallelism for SPHinXsys

Abstract
Simulations based on particle methods, such as Smoothed Particle Hydrodynamics (SPH), are known to be computationally demanding. While such methods have for long been executed in parallel on multi-core CPUs, in recent years the increasing adoption of many-core accelerators, such as GPUs. However, hardware fragmentation and vendor-specific programming interfaces are still characterizing their market. Hence, support for various hardware configurations may easily lead to non-trivial and less maintainable implementations. To leverage over some higher-level specifications have become available recently, such as the SYCL programming standard, this work highlights the initial effort in adopting the SYCL standard for the execution of SPHinXsys, an open-source multi-physics library. The result is an execution model able to run the same implementation on variable (heterogeneous) hardware, with considerable speed-up compared to the current multi-core CPU parallelization. Among others, representation of data-structures for parallel access, communication strategies, and parallel methods for data sorting will be topics discussed in depth. Benchmarks has also been presented, showcasing performance comparisons between the current multi-core CPU implementation and the newly introduced SYCL parallelization with a GPU back-end.
keywords:
Smoothed particle hydrodynamics , GPU acceleration , SYCL , data-structure
journal: Elsevier
1 Introduction
As smoothed particle hydrodynamics (SPH) is a typical particle-based method, many SPH libraries suffer from high computational costs due to intensive particle interactions. Compared to the multi-cores CPU system, the many-cores device (typically GPU) system provides a much higher performance/cost ratio for the intermediate (multi-million particles) scale simulations, which are frequently encountered in industrial applications. On the other hand, since the versatile functionality of CPU system, it is expected that SPH libraries are using heterogeneous parallelism so that they are able to take advantage of the both.
Currently, some SPH libraries already offer GPU support, either through vendor-specific implementations [1] [2] [3], or heterogeneous programming using the OpenCL standard [4] [5]. An important issue of these implementations is that they require computing kernels to be defined, as a low-level specification, with specific directives separated from the rest of the code, rendering it less easily programmable and maintainable.
2 Open-source multi-physics library SPHinXsys
SPHinXsys [6], is an open-source C++ SPH multi-physics simulations library. It addresses the complexities of fluid dynamics, structural mechanics, fluid-structure interactions, thermal analysis, chemical reactions and AI-aware optimizations. SPHinXsys has offered CPU parallelism using Intel’s Threading Building Blocks (TBB) library [7] and several features suitable for open-source based development.
First, since the parallel execution is encapsulated into the low-level classes decoupled from the SPH method, the developers, as typical mechanical engineers, only need to work the numerical discretization without concerning parallelization. Second, variable testing approaches, including unit test, google test and regression test, have been incorporated so that refraction, adding new features and other modifications can be implemented without the worries about the already established functionalities. Third, automated cross-platform continuous integration/development (CI/CD) tests have being carried out on the open-source development platform frequently for any modification of the main branch of SPHinXsys repository. In addition, SPHinXsys makes the uses of several third particle libraries, such as Simbody library for multi-body dynamics, Pybind11 library for generating python interface used for machine learning and optimization applications.
3 SYCL Programming Standard
Different from the other GPU-able SPH libraries, our work is based on the SYCL standard [8] for parallel computing. SYCL specification defines a new single-source ISO C++ compliant standard with compute acceleration. It aims to be an high level abstraction that can be programmed as classical CPU code, without accelerator-specific directive and application interface (API) calls. In this work, the Intel SYCL implementation, i.e. Data Parallel C++ (DPC++) [7] is chosen, because it is a part of Intel’s oneAPI suit which also includes the TBB library already used in SPHinXsys.
A SYCL computing kernel (sharing the same term with the smoothing kernel used in SPH) is a scoped block of code executed on a SYCL device under the global context queue and the local one command_group initiated by the host. Note that, as the SYCL host and device both can be CPU, CI/CD jobs can test SYCL kernels on computers without GPU, such the standard runner provided by Github platform.
4 Implementing SYCL kernels in SPHinXsys
The long term aim for SPHinXsys development is employing unified codebase for the SPH methods so the multi-physics simulation can be carried on heterogeneous computing systems in a collaborative way. Therefore each SPH algorithm can be executed on host or device in sequential or parallel.
Before implementing SYCL kernels, SPHinXsys already has defined two execution policies, namely sequenced_policy and parallel_policy for CPU computing. We deliberately implement the sequential execution on CPU because it is essential for easy debug propose. As SPHinXsys splits the execution, namely particle_for and particle_reduce, and SPH methods, i.e. the classes inherited from the base local_dynamics with the same source code for both executions.
Our basic concept on implementing SYCL kernels is to extend the same programming pattern for CPU computing by adding an extra execution policy, namely parallel_device_policy to identify that the same source code of SPH methods will be executed in parallel on the device.
Though the concept is straightforward, its implementation is not trivial since it relies on the an important assumption that the codes written in computing kernels should work for all execution policies. However, the source code of SPH methods previously used for CPU computing generally are not executable on device due to the widely used references, virtual functions and standard library dynamic memory allocations. Therefore, a extended program pattern based on SYCL unified share memory (USM) is developed. The extended pattern has three layer, i.e. outer, kernel-shell and kernel, dependent on their distances to computing kernel which is the only one executed by the execution algorithms.
The outer layer is the same as before and defines the physical problem. Within this layer, the global variables are defined, the preprocess is done by generating particles respect to geometric information. Third-parties environment and methods, e.g. Simbody environment and methods, are referred directly. The kernel-shell layer is the SPH method definition layer, which can be obtained from the CPU computing code with minimum modification. This layer is used to define individual SPH method. At this layer, the interface data structures, namely discrete_ and singular_variables are used to transfer the outer-layer data to those used in computing kernel. The kernel layer defines all the computing via kernels, and is obtained by splitting out the computing function from original method classes. The objects of this layer are only instantiated and dispatched at the beginning of computing with the help of the kernel-shell layer interface and execution implementation. Note that the data transformed from kernel-shell layer has the form of raw pointer, and can be accessed on host and device just as arrays.
5 Cell linked-list, direct search and sorting
Since the limitation of device hardware, SYCL specification does not allow the usage of concurrent_vector, which is has been used in SPHinXys for constructing the cell linked-list for particle neighbor searching algorithms. To achieve high performance on constructing cell linked-list on device concurrently, atomic operations are used to avoid the possible thread conflicts.
Again, due to the memory limit of GPU, the full particle configuration, including neighbor list, kernel values used in SPhinXsys can not be used anymore. Therefore, direct search is used, i.e. the kernel values will be computed for all particles within the nearest cells in the computing kernels directly.
Similarly, due to poor performance of GPU on recursive algorithms, the particle sorting executed by device uses radix_sort other than quick_sort. Note that, this is the only computing kernel that used different codes for device execution.
6 Performance Evaluation
The dam-break flow simulation which very often has been used as reference is chosen as the benchmark test-case. Additionally, simulations have been executed with single and double floating-point precision. Computations are carried out on a workstation composed of two Intel Xeon E5-2603 v4 and one NVIDIA GeForce RTX 2080 Ti. Results presented in Figs.1 and 2 delineate a performance improvement up to 27 times the execution on CPU using single precision.
Figure 1: Simulation of 2-dimensional simulation of dam-break flow.
Figure 2: Simulation of 3-dimensional simulation of dam-break flow.
In particular, it is clear how single floating-point precision on GPU runs considerably better than any other configuration when executed using direct search.
In order to compare the achieved results with an existing GPU parallelization of SPH, SPHinXsys has been tested against DualSPHysics, an SPH solver parallelized with CUDA. The test-case defined by Kleefsman et al. in [9] is used to compare the two implementations. The baseline results of DualSPHysics have been taken from an existing benchmark [10] involving the same test-case. In particular, the results considered are simulation executed on an NVIDIA RTX 2080Ti. The same test-case has been replicated in SPHinXsys and simulated on the same GPU. As shown in Fig. 3, the present simulation results are in quite good agreement with the literature.
Figure 3: Snapshots for the dam-break with obstacle test-case.
Figure 4: Runtime comparison between DualSPHysics and SPHinXsys for the dam-break with obstacle test-case.
Figure 5: Comparison between DualSPHysics and SPHinXsys based on the number of particles interacting each second.
Results are presented in Figure 4, and they show a SPHinXsys runtime that is half of the one needs by DualSPHysics. Specifically, for the case with the maximum number of particles of 56 million, while DualSPHysics is able to computed the first two seconds of the simulation in 32 hours, SPHinXsys simulates the same period in 18 hours. To confirm such results, Giga Particle Interactions Per Second (GPIPS) has also been computed, which is considered by Domínguez et al. [1] a better unit of measurement compared to pure runtime. The latter could in fact be affected by different timestep sizes, I/O output, etc. It is found that SPHinXsys GPIPS are twice the amount of DualSPHysics, effectively confirming the runtime results, as shown in Figure 5.
7 Conclusions
In this work, we presents the realization of heterogeneous parallelism for SPHinXsys based on the SYCL standard for GPU computing. This is motivated by the expectation that, besides accelerating the SPH simulation and easy programming and maintaining, SPHinXsys is still able to maintain the already offered benefits for open-source based development without notable compensation.
As the present implementation are actually more on generalizing the usage of data types so that single computing kernel can be executed on host and device, sequenced or parallelized, other than the exact numerical algorithm itself. Heterogeneous parallelism is achieved with the minimum modification of the original program pattern. This is especially suitable for the engineer developer as they do not need understand execution details of the numerical method they are developing.
Another advantage is that, since generally only a single source code for the computing kernel to be executed with all execution policies, SPHinXsys allows for the development and testing of numerical methods even in environments without GPUs or even DPC++ installed. If the source code for a computing kernel is crafted following our specified guidelines and prove functional in a standard platform, e.g. Linux system using the GNU compiler, they will seamlessly operate in environments equipped with DPC++ and GPU support.